# From Moderation to Mediation: Can LLMs Serve as Mediators in Online Flame Wars?

### v2

@@ 4 Data Collection
Target-user Identification and Thread Extraction. We again used Gemini-2.5 to detect the two users with the most flame-war interactions in their comments. After identifying the two users’ names, we initially extracted only the direct replies exchanged between them. However, we observed that many relevant subthreads included additional participants whose replies contributed to the escalation or resolution of the conflict. To preserve the full conversational context, we expanded our extraction to include all nested replies within any subtree where both target users appeared, thereby capturing the dynamics of the surrounding dialogue while maintaining a focus on their exchange. After applying this filtering, the final dataset comprised 504 posts with rich conversational structures suitable for mediation generation and principle-based evaluation. Table 1 presents the detailed statistics of our collected dataset.
@@ 5.2 User Simulation Result
Generation of Intervened Conversations. We construct a user simulator that models how conflicting users in online discussions might respond to mediator interventions. Given an original discussion thread $G$, the simulator generates an intervened thread $G^{\prime}$ by inserting a model-generated mediation output $J(G)$ or $S(G)$ at an appropriate turn. The simulator then predicts the subsequent user responses conditioned on both $G$ and J(G)J(G)/$S(G)$, thereby approximating post-intervention dynamics. This setup enables an evaluation of whether mediation messages can successfully reduce hostility, encourage cooperative tone shifts, or promote topic re-alignment in extended interactions.

### v5

# arXiv:2512.03005v5[cs.AI] 26 Feb 2026

@@ 4 Data Collection
Target-user Identification and Thread Extraction. We again used Gemini-2.5 to detect the two users with the most flame-war interactions in their comments. After identifying the two users’ names, we initially extracted only the direct replies exchanged between them. However, we observed that many relevant subthreads included additional participants whose replies contributed to the escalation or resolution of the conflict. To preserve the full conversational context, we expanded our extraction to include all nested replies within any subtree where both target users appeared, thereby capturing the dynamics of the surrounding dialogue while maintaining a focus on their exchange. After applying this filtering, the final dataset comprised 737 posts with rich conversational structures suitable for mediation generation and principle-based evaluation. Table 1 presents the detailed statistics of our collected dataset.
@@ 5.2 User Simulation Result
Generation of Intervened Conversations. We construct a user simulator that models how conflicting users in online discussions might respond to mediator interventions. Given an original discussion thread $G$, the simulator (here we use Qwen3-4B as the simulator model) generates an intervened thread $G^{\prime}$ by inserting a model-generated mediation output $J(G)$ or $S(G)$ at an appropriate turn. The simulator then predicts the subsequent user responses conditioned on both $G$ and J(G)J(G)/$S(G)$, thereby approximating post-intervention dynamics. This setup enables an evaluation of whether mediation messages can successfully reduce hostility, encourage cooperative tone shifts, or promote topic re-alignment in extended interactions.
@@ 7 Conclusion
Acknowledgement
This work was supported by National Science Foundation Awards No. 2435164 and 2435165, and a Google Award for Inclusion Research. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing official policies, either expressed or implied, of NSF.